# Existence, uniqueness, and regularity of

### v1

@@ Abstract
We establish the existence, uniqueness, and $W^{1,2,p}$-regularity of solutions to fully nonlinear parabolic obstacle problems when the obstacle is the pointwise supremum of arbitrary functions in $W^{1,2,p}$ and the operator is only assumed to be measurable in the state and time variables. The results hold for a large class of non-smooth obstacles, including all convex obstacles. Applied to stopping problems, they imply that the decision maker never stops at a convex kink of the stopping payoff. The proof relies on new $W^{1,2,p}$-estimates for obstacle problems where the obstacle is the maximum of finitely many functions in $W^{1,2,p}$.
1 Introduction and Main Result
We study the following fully nonlinear parabolic obstacle problem with Dirichlet boundary data on the domain $\mathcal{Y}=[0,T)\times\mathcal{X}$ where $T$ is finite, $d\in\mathbb{N}$, and $\mathcal{X}$ is a bounded, open subset of $\mathbb{R}^{d}$:
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ in }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}\end{cases}$ (1)
@@ 1 Introduction and Main Result
We are interested in obstacles $g$ that take the form $g=\underset{a\in A}{\sup}\,g^{a}$ for functions that lie in the Sobolev space $W^{1,2,p}\left(\mathcal{Y}\right)$ for some $p\in(d+2,\infty]$. This form arises in stopping problems for which the decision maker must decide, upon stopping, on some action $a$. In this case, $g^{a}(t,x)$ is the stopping payoff when stopping at time $t$ in state $x$ and taking action $a$ upon stopping. A decision maker who stops at time $t$ would then optimally choose the action $a$ that maximizes $g^{a}(t,x)$ over all possible actions $a\in A$, and the resulting stopping payoff, i.e., obstacle, is $g$.
Such problems are pervasive, for example, in information acquisition models where “stopping” means concluding the information acquisition stage and $a$ is the decision taken after that stage. Examples include Wald, (1992), Décamps et al., (2006), Fudenberg et al., (2018) and Camboni and Durandard, (2024). In these and other instances, knowing that the Hamilton-Jacobi-Bellman equation has a solution proves instrumental in deriving the optimal policy and can lead to important insights (Dixit,, 1993).
While commonly arising in applications, these problems create a challenge because the obstacle $g$ is irregular. In particular, $g$ will typically have kinks whenever the optimal action $a$ changes as a function of $x$ and $A$ is finite.
Our main result requires the following conditions (all notation and definitions are formally introduced in Section 2).
@@
This assumption allows domains $\mathcal{X}$ with non-smooth boundaries as long as they are Lipschitz. Informally, $\mathcal{X}$ is Lipschitz if its boundary $\partial\mathcal{X}$ can be viewed locally as the graph of a Lipschitz function for some coordinate system. Allowing non-smooth domains is particularly useful for economic and financial applications, whose domains are often given by a positive orthant (e.g., in the case of price vectors), a unit cube, or a simplex (e.g., when $x$ is a probability distribution representing the decision maker’s belief about a state of the world with $d$ possible values). All these domains fail to be smooth, but they all satisfy the Lipschitz property.
The next two assumptions concern the primitives $b$ and $g$.
@@
$g=\underset{a\in A}{\sup}\;g^{a}$ where $A$ is a separable topological space and (i) $g^{a}\in W^{1,2,p}(\mathcal{Y})$ for all $a$ and $\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}<\infty$, (ii) $g^{a}\leq b$ on $\partial\mathcal{Y}$, and (iii) viewed as a map from $A$ to $C^{0}\left(\mathcal{Y}\right)$, the map $a\mapsto g^{a}$ is continuous.
Part (iii) of Assumption 3 guarantees that the obstacle can be approximated by the maximum of finitely many $W^{1,2,p}$-functions, and hence the validity of the approximation argument in the proof of Theorem 1. It is a sufficient condition: The proof only requires the existence of an approximating sequence $(g^{n}=\underset{a\in A_{n}}{\max}\,g^{a})_{n}$, with $A_{n}$ finite for all $n$.
@@
$\begin{cases}&\mathcal{P}_{\lambda,\Lambda}^{-}\left(M-\tilde{M}\right)-\omega_{1}\left(\right|r-\tilde{r}\left|\right)-\omega_{2}\left(\right|q-\tilde{q}\left|\right)\\ &\quad\quad\leq F\left(t,x,r,q,M\right)-F\left(t,x,\tilde{r},\tilde{q},\tilde{M}\right)\\ &\quad\quad\quad\quad\leq\mathcal{P}_{\lambda,\Lambda}^{+}\left(M-\tilde{M}\right)+\omega_{1}\left(\right|r-\tilde{r}\left|\right)+\omega_{2}\left(\right|p-\tilde{p}\left|\right).\end{cases}$ (SC)
@@
This assumption implies in particular the following properties: (i) $F$ is uniformly elliptic, (ii) $F(t,x,r,p,M)$ is continuous in $r$, $p$, and $M$ uniformly in $(t,x)$, and (iii) $F$ grows at most linearly in $(t,x)$.
Assumption 4 is weaker than the conditions typically imposed on $F$ in the literature: First, $F$ is not required to be Lipschitz in $(r,q)$. Second, $F$ need not be continuous in $(t,x)$. Third, $F$ need not be monotonic in $r$ [^2] We do impose a weaker monotonicity condition (Assumption 8), for our uniqueness result.
@@
$F(t,x,0,0,M)$ is convex in $M$ for all $(t,x)$. [^3] The concave case follows (upon modifying Assumption 8 appropriately), since $u$ solves $\max\{g-v,F(t,x,v,v_{x},v_{xx})+v_{t}\}=0$ if and only if it solves $\min\{v-g,-F(t,x,v,v_{x},v_{xx})-v_{t}\}=0$.
@@
As pointed out in Dong, (2020), Assumption 6 is the weakest known assumption under which a solution in $W^{1,2,p}$ always exists for Dirichlet problems, even in the linear case. Moreover, as the counter-examples in Meyers, (1963) or ural1967невозможности suggest, it may be close to optimal.
@@
The final assumption is used to apply a comparison principle to (1). It is not needed for our existence results, but is essential to guarantee uniqueness of the solution to (1).
@@
There exists $\kappa>0$ such that, for all $(t,x)\in\mathcal{Y}$, $p\in\mathbb{R}^{d}$, and $M\in\mathbb{S}^{d}$,
$r\to F(t,x,r,p,M)-\kappa r$
@@
Our main result [^4] The structure condition (SC) guarantees that any solution of $u_{t}+F[u]=0$ with $u=b$ on $\partial\mathcal{Y}$ is bounded below. Therefore, by choosing $g$ small enough, Theorem 1 also guarantees the existence, uniqueness, and regularity of the solution without obstacles. establishes the existence, uniqueness, and regularity for solutions to (1).
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 17 hold. Then (1) has an $L^{p}$-solution $u$.
Moreover, there exists $C=C\left(d,p,\lambda,\Lambda,R,diam\left(\mathcal{X}\right),T,dist(\mathcal{Y}^{\prime},\mathcal{Y})\right)\in\mathbb{R}_{+}$ such that, for all compact $\mathcal{Y}^{\prime}\subset\mathcal{Y}$,
$\left\|u\right\|_{W^{1,2,p}\left(\mathcal{Y^{\prime}}\right)}\leq C\left(1+\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right),$ (2)
and a modulus of continuity $\bar{\omega}$ depending only on $p,d,\lambda,\Lambda,R,T,diam\left(\mathcal{X}\right),\underset{a\in A}{\sup\,}\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}$, $\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)},\left\|b\right\|_{L^{\infty}\left(\partial\mathcal{Y}\right)}$, the modulus of continuity of $b$, and the parameters of the cone condition of $\mathcal{X}$, such that, for all $(t,x),(t^{\prime},x^{\prime})\in\bar{\mathcal{Y}}$,
$\left|u(t,x)-u(t^{\prime},x^{\prime})\right|\leq\bar{\omega}\left(\left|(t,x)-(t^{\prime},x^{\prime})\right|\right).$

### v2

@@ Abstract
We establish the existence, uniqueness, and $W^{1,2,p}$-regularity of solutions to fully-nonlinear, parabolic obstacle problems when the obstacle is the pointwise supremum of functions in $W^{1,2,p}$ and the nonlinear operator is required only to be measurable in the state and time variables. In particular, the results hold for all convex obstacles. Applied to stopping problems, they provide general conditions under which a decision maker never stops at a convex kink of the stopping payoff. The proof relies on new $W^{1,2,p}$-estimates for obstacle problems when the obstacle is the maximum of finitely many functions in $W^{1,2,p}$.
1 Introduction
We study a fully nonlinear parabolic obstacle problem with Dirichlet boundary data on the domain $\mathcal{Y}=[0,T)\times\mathcal{X}$ where $T$ is finite and $\mathcal{X}$ is a bounded, open subset of $\mathbb{R}^{d}$ for some $d\in\mathbb{N}$:
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ on }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (1)
@@ 1 Introduction
Parabolic obstacle problems consist of finding the smallest function that (i) exceeds a given obstacle function and (ii) is a supersolution of a given parabolic equation (Petrosyan and Shahgholian,, 2007). Such a function solves an equation of the form (1) and is called the solution of the obstacle problem. Obstacle problems arise in physics, e.g., to study phase transitions (Lamé and Clapeyron, (1831); Stefan, (1889); Friedman, (1982)), in biology, e.g., to study tumor growth (Greenspan, (1976); Bazaliy and Friedman, (2003)), in economics, e.g., to study learning and investment decisions (Wald, (1947); Dixit, (1993); Moscarini and Smith, (2001); Décamps et al., (2024)), and in finance, e.g., to study American options (Jacka, (1991); Shiryaev, (1999); Villeneuve, (1999)).
We are particularly interested in obstacle problems that arise in optimal stochastic control and stopping problems, whether they involve a single decision maker or multiple players interacting in a stochastic game. When the primitives of the corresponding obstacle problem are sufficiently regular, it is well-known that the value function for the optimal control and stopping of a diffusion is the solution of an obstacle problem where the operator $F$ is the Hamilton-Jacobi-Bellman operator and the obstacle is the stopping payoff (Bensoussan and Lions,, 1978; Friedman,, 1982; Karatzas and Sudderth,, 2001; Peskir and Shiryaev,, 2006; Petrosyan et al.,, 2012; Strulovici and Szydlowski,, 2015).
In many applications, however, the operator, the obstacle, or the domain that define the obstacle problem fail to be regular. In stochastic games, for example, the Hamilton-Jacobi-Bellman operator that defines a player’s optimization problem has coefficients that depend on the strategies of other players and may for this reason fail to be continuous (Kuvalekar and Lipnowski,, 2020). Indeed, such discontinuities typically arise when players choose at each instant from finite action sets.
@@
Similarly, the obstacle $g$ that arises in many stopping problems takes the form $g=\underset{a\in A}{\sup}\,g^{a}$, and may thus have kinks even when the individual functions $g^{a}$ are smooth. This form arises when the decision maker must decide, upon stopping, on some action $a$. In this case, $g^{a}(t,x)$ is the stopping payoff when stopping at time $t$ in state $x$ and taking action $a$ upon stopping. A decision maker who stops at time $t$ would then optimally choose the action $a$ that maximizes $g^{a}(t,x)$ over all possible actions $a\in A$, and the resulting stopping payoff, i.e., obstacle, is $g$. Such problems are pervasive in information acquisition models where “stopping” means concluding the information acquisition stage and $a$ is the decision taken after that stage (Wald, (1992), Décamps et al., (2006), Fudenberg et al., (2018) and Camboni and Durandard, (2024)).
Finally, in many obstacle problems that arise in economics and other fields, the domain of the state variable fails to be smooth. In information acquisition problems, for example, the state variable is a belief and the relevant domain is a probability simplex, which has kinks at its vertices (and, more generally, along the lower-dimensional faces of the probability simplex). As another example, the state variable in many finance and general equilibrium problems is a price vector whose relevant domain is a positive orthant, which again fails to be smooth.
To address the challenges created by a lack of regularity in the primitives of obstacle problems, researchers have often used techniques tailored to their specific application. For instance, Décamps et al., (2006) consider the elliptic obstacle equation associated with an optimal stopping problem when the payoff upon stopping (i.e., the obstacle) is the maximum of two smooth convex functions, and use a probabilistic local time argument to show that the decision-maker never stops at the convex kink. The results for regular obstacles then apply, and the authors can show that the value function is a smooth solution of (1).
This paper aims to address these challenges all at once. Our main result, Theorem 1, establishes the existence, uniqueness, and $W^{1,2,p}$-regularity of a solution to fully nonlinear obstacle problems when (i) the operator is required only to be measurable in $(t,x)$, (ii) the obstacle is the supremum of functions in $W^{1,2,p}$—and may thus have kinks—and (iii) the domain is required only to be Lipschitz—and may thus have corners and other kinks. Theorem 1 covers various applications considered in economics. For instance, it allows us to recover Décamps et al., (2006)’s result when combined with a Sobolev embedding theorem. [^2] Corollary 1 makes this point formally. More generally, it guarantees that the solution of (1) is continuously differentiable in space and, thus, that the decision-maker never stops at a point of nondifferentiability of the stopping payoff $g$. Since the results apply to any convex obstacle, they imply that any point of nondifferentiability of the stopping payoff must be in the continuation region.
Results similar to Theorem 1 have appeared under stronger assumptions. A precursor is Friedman, (1982), which shows that when the primitives are smooth and the operator $F$ is linear, the obstacle problem has a unique solution in $W^{1,2,\infty}\left({\mathcal{Y}}\right)$. Petrosyan and Shahgholian, (2007); Audrito and Kukuljan, (2023) extend this result to nonlinear operators when the obstacle and operator are smooth and the operator satisfies strong convexity and growth conditions. When the operator is not smooth, solvability and regularity results for elliptic and parabolic equations in the absence of obstacles appear in Caffarelli, (1989), Escauriaza, (1993), Caffarelli et al., (1996), Winter, (2009), and Krylov, (2010) for the elliptic case, and in Crandall et al., (1998), Crandall et al., (1999), Crandall et al., (2000), Dong et al., (2013), and Krylov, (2017) for the parabolic case, among others. The recent monograph Krylov, (2018) offers an up-to-date general treatment. Finally, Byun et al., (2018, 2022) study the regularity of viscosity solutions for obstacle problems when the obstacle is smooth and one of the following conditions hold: the operator is linear or the problem is elliptic.
We now describe more specifically how our main theorem generalizes existing results. Crandall et al., (2000) and Krylov, (2017, 2018) prove the existence, uniqueness, and $W^{1,2,p}$-regularity of solutions of fully nonlinear parabolic problems in the absence of an obstacle (let alone an irregular one). In addition, Krylov, (2017, 2018) assume that the domain is smooth and Crandall et al., (2000) assume that the operator is Lipschitz continuous in the derivatives of the value function, while the theorem of the present paper requires only continuity. [^3] Krylov, (2017, 2018) use the same concept of solutions as we do: $L^{p}$-solutions, while Crandall et al., (2000) work with the weaker concept of $L^{p}$-viscosity solutions. However, these concepts coincide under our assumptions on the nonlinear operator. Byun et al., (2018) proposes a new method to prove that parabolic obstacle problems on smooth domains are solvable in $W^{1,2,p}$ when the operator is linear and the obstacle is in $W^{1,2,p}$, a condition that is not satisfied by the kinked obstacles that arise in various stopping problems. Finally, Byun et al., (2022) adapt the technique in Byun et al., (2018) to study fully nonlinear elliptic obstacle problems in smooth domains when the obstacle is in $W^{2,p}$.
Theorem 1 builds on the results and ideas in these earlier papers to obtain the existence, uniqueness, and $W^{1,2,p}$-regularity of the solution to fully nonlinear parabolic obstacle problems when the operator is measurable in $(t,x)$, the obstacle is irregular, and the domain is not required to be smooth. The proof of Theorem 1 relies on PDE methods and generalizes existing results derived by probabilistic methods in the context of optimal stopping. Specifically, we extend the approximation argument developed in Byun et al., (2018) and Byun et al., (2022) to the fully nonlinear parabolic case and allow non-smooth obstacles. To do so, we first obtain a $W^{1,2,p}$-estimate when the obstacle is in $W^{1,2,p}$ and the domain is smooth using known estimates on the regularity of solutions of the Dirichlet problem, e.g., Theorem 12.1.7 and 15.1.3 in Krylov, (2018). We then show by induction that the existence and regularity results for a single $W^{1,2,p}$-obstacle carry to the pointwise maximum of a finite number of $W^{1,2,p}$-obstacles (Lemma 2). Finally, we extend the result to Lipschitz domains and obstacles that are the supremum of a separable family of functions in $W^{1,2,p}$ by a limit argument.
1.1 Outline of the Paper
Section 2 introduces notations and definitions. Section 3 formally states our assumptions and main result, Theorem 1. Section 4 contains two results needed for the proof of Theorem 1: (i) a comparison principle that applies to fully nonlinear parabolic obstacle problems with measurable ingredients (Proposition 1), and (ii) a stability theorem for solutions to obstacle problems in $W^{1,2,p}\left(\mathcal{Y}\right)$ (Theorem 2). Section 5 establishes our main result in a simpler setting in which the obstacle is the maximum of finitely many functions in $W^{1,2,p}$ and the domain $\mathcal{X}$ is smooth. Finally, Section 6 proves Theorem 1 in full generality by first extending the result of Section 5 to general separable sets $A$ and then extending this result further to non-smooth domains.
2 Preliminaries
2.1 Notation
  • For $D>0$, $B_{D}(x)$ denotes the open ball of radius $D$ centered around $x$ and $C_{D}(t,x)=[t,t+D)\times B_{D}(x)$.
  • $\mathcal{C}^{k}\left(\mathcal{Z}\right)$ is the space of continuous functions on $\mathcal{Z}$ if $k=0$ and of $k$-times continuously differentiable functions on $\mathcal{Z}$ if $k\geq 1$.
  • $W^{1,2,p}\left(\mathcal{Z}\right)$ denotes the Sobolev space of functions defined on the set $\mathcal{Z}$ whose first weak time derivative and second weak space derivatives are $L^{p}$-integrable. $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ is the space of functions that belong to $W^{1,2,p}\left(\mathcal{Y}^{\prime}\right)$ for all compact subsets $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$.
  • For a function $u\in W^{1,2,p}_{loc}\left(\mathcal{Z}\right)$, $u_{t},u_{x}$, and $u_{xx}$ stand for the first weak derivatives of $u$ with respect to $t$, the first weak derivatives of $u$ with respect to $x$, and the second weak derivatives of $u$ with respect to $x$, respectively.
[…]
  • For $k\in\mathbb{N}$, an open bounded subset $\mathcal{Z}$ of $\mathbb{R}^{d}$ is $\mathcal{C}^{k,Lip}$ if, for every $x\in\partial\mathcal{Z}$, there exists a neighborhood $V$ of $x$, $\alpha>0,\beta>0$, an affine map $T:\mathbb{R}^{d-1}\times\mathbb{R}\to\mathbb{R}^{d}$, and a map $\phi:B_{\alpha}(0)\subseteq\mathbb{R}^{d-1}\to\mathbb{R}$ that is Lipschitz continuous if $k=0$ or $k$-times continuously differentiable with Lipschitz derivatives if $k\geq 1$, such that
    $\partial{\mathcal{Z}}\cap V=T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\eta=\phi(\xi)\right\}\right),$
    and
    $T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\phi(\xi)<\eta<\phi(\xi)+\beta\right\}\right)\subseteq{\mathcal{Z}}\cap V$
    and $T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\phi(\xi)-\beta<\eta<\phi(\xi)\right\}\right)\subseteq\mathbb{R}^{d}\setminus\left(\bar{{\mathcal{Z}}}\cap\bar{V}\right).$
    The first part of the definition requires that there locally exist a coordinate system such that the boundary coincides locally with the graph of a function whose $k^{th}$ derivative (or the function itself if $k=0$) is Lipschitz. The two inclusions in the second part of the definition guarantee that the interior and the exterior of the domain are nonempty and locally contain a cone around any point of the boundary. They rule out, for example, the domain $\{(x,y)\,:\,x>0,\,|y|<x^{2}\}\subset\mathbb{R}^{2}$. Informally, a domain $\mathcal{Z}$ is Lipschitz if its boundary $\partial\mathcal{Z}$ can be viewed locally as the graph of a Lipschitz function for some coordinate system.
  • The $\mathcal{C}^{k,Lip}$-norm of the boundary of a $\mathcal{C}^{k,Lip}$ open bounded subset $\mathcal{Z}$ is defined by $\left\|\partial\mathcal{Z}\right\|_{\mathcal{C}^{k,Lip}}=\underset{x\in\partial\mathcal{Z}}{\sup}\,\left\|\phi\right\|_{\mathcal{C}^{k,1}(V)}$ where $V$ and $\phi$ depend on $x$ and are as defined above. Note that the $\mathcal{C}^{k,Lip}$-norm of the boundary of a bounded $\mathcal{C}^{k,Lip}$-set in $\mathbb{R}^{d}$ is finite, because in this case the boundary is compact.
  • We will often use the letter $C$ as a generic letter to denote different bounds appearing in various estimates, and which can be explicitly computed in terms of primitives of the problem.
2.2 Solution concepts and definitions
Definition 1
A function $u$ is an $L^{p}$-subsolution (respectively, $L^{p}$-supersolution) of (1) if the following conditions hold: (i) $u\in{W}^{1,2,p}_{loc}(\mathcal{Y})\cap\mathcal{C}^{0}(\bar{\mathcal{Y}})$, (ii) $u\leq b$ (respectively, $\geq b$) on $\partial{\mathcal{Y}}$, and (iii)
$\max\left\{\,u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\,\,(\text{respectively, }\leq 0)\text{ a.e. on }\mathcal{Y}.$
We will also use viscosity solutions in stating and applying the comparison principle of Section 4.1. There are several concepts of viscosity solutions depending on the set of test functions used in the definition. We use test functions in $W^{1,2,p}\left(\mathcal{Y}\right)$, which corresponds to what is sometimes called “$L^{p}$-viscosity solutions.” [^4] When $F$ is continuous, one can take the test functions to be in $\mathcal{C}^{1,2}$. In this case, the solution is called a $\mathcal{C}$-viscosity solution. In this case, the concepts of $\mathcal{C}$-viscosity solutions and of viscosity solutions as we defined them coincide (Lemma 2.9 in Crandall et al., (2000)).
Definition 2
A function $u:\bar{\mathcal{Y}}\to\mathbb{R}$ is a viscosity subsolution (respectively, supersolution) of (1) if (i) it is continuous, (ii) $u\leq b$ (respectively, $\geq b$) on $\partial{\mathcal{Y}}_{T}$, and (iii) for all $(t_{0},x_{0})\in\mathcal{Y}$ and all $\varphi\in{W}^{1,2,p}_{loc}(\mathcal{Y})$ such that $u-\varphi$ (respectively $\varphi-u$) has a maximum at $(t_{0},x_{0})$, one has
$\underset{(t,x)\to(t_{0},x_{0})}{\mathrm{ess}\limsup}$ $\,\max\left\{\,\varphi_{t}+F\left(t,x,u,\varphi_{x},\varphi_{xx}\right),\,g-u\right\}\geq 0$
[…]
Lemma 1
Let $u$ be an $L^{p}$-subsolution (respectively, supersolution) of (1). Then $u$ is a viscosity subsolution (respectively, supersolution) of (1).
3 Main Result
Our main result is derived under the following assumptions.
@@ 3 Main Result
This assumption allows domains $\mathcal{X}$ with non-smooth boundaries. As noted in the Introduction, allowing for such domains is particularly useful in economic and financial applications, where domains are often equal to a positive orthant (e.g., in the case of price vectors), a unit cube, or a simplex (e.g., when $x$ is a probability distribution representing the decision maker’s belief about a state of the world with $d$ possible values). These domains all fail to be smooth, but they all satisfy the Lipschitz property.
The next two assumptions concern the exogenous boundary payoff function $b$ and the stopping payoff function $g$.
@@
Assumption 2 implies that $b$ is bounded on $\bar{\mathcal{Y}}$, a property that we will use when computing estimates.
@@
$g=\underset{a\in A}{\sup}\;g^{a}$ where $A$ is a separable topological space and the functions $g^{a}$ have the following properties: (i) $g^{a}\in W^{1,2,p}(\mathcal{Y})$ for all $a$ and $\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}<\infty$, (ii) $g^{a}\leq b$ on $\partial\mathcal{Y}$, and (iii) the map $a\mapsto g^{a}$ is continuous from $A$ to $C^{0}\left(\mathcal{Y}\right)$.
Part (iii) of Assumption 3 guarantees that the obstacle can be approximated by the maximum of finitely many $W^{1,2,p}$-functions, and allows us to use an approximation argument in the proof of Theorem 1.555Part (iii) is sufficient but need not be necessary. The proof requires only the existence of an approximating sequence $(g^{n}=\underset{a\in A_{n}}{\max}\,g^{a})_{n}$, with $A_{n}$ finite for all $n$.
@@
$\begin{cases}&\mathcal{P}_{\lambda,\Lambda}^{-}\left(M-\tilde{M}\right)-\omega_{1}\left(\right|r-\tilde{r}\left|\right)-\omega_{2}\left(\right|q-\tilde{q}\left|\right)\\ &\quad\quad\leq F\left(t,x,r,q,M\right)-F\left(t,x,\tilde{r},\tilde{q},\tilde{M}\right)\\ &\quad\quad\quad\quad\leq\mathcal{P}_{\lambda,\Lambda}^{+}\left(M-\tilde{M}\right)+\omega_{1}\left(\right|r-\tilde{r}\left|\right)+\omega_{2}\left(\right|q-\tilde{q}\left|\right).\end{cases}$ (SC)
@@
Assumption 4 implies that $F$ is uniformly elliptic, that $F(t,x,r,p,M)$ is continuous in $r$, $p$, and $M$ uniformly over $(t,x)$, and that $F$ grows at most linearly in the value and first derivatives of the value function.
The assumption is weaker than the assumptions typically imposed on $F$ in the literature for the following reasons: (i) $F$ is not required to be Lipschitz in $(r,q)$; (ii) $F$ need not be continuous in $(t,x)$; (iii) $F$ need not be monotonic in $r$ [^6] To establish the uniqueness part of our result, and only this part, we do impose a weak monotonicity condition (Assumption 8).
@@
$F(t,x,0,0,M)$ is convex in $M$ for all $(t,x)$. [^7] When $F$ is concave case, our results remain valid (upon modifying Assumption 8 appropriately), since $u$ solves $\max\{g-v,F(t,x,v,v_{x},v_{xx})+v_{t}\}=0$ if and only if it solves $\min\{v-g,-F(t,x,v,v_{x},v_{xx})-v_{t}\}=0$.
@@
Assumption 6 is needed to ensure that the weak derivatives of solutions to canonical equations related to obstacle problem exist a.e. and belong to $L^{p}$, by guaranteeing that the coefficient of the highest-order derivative in the linearized version of the operator $F$ is regular enough in $(t,x)$. As pointed out in Dong, (2020), Assumption 6 is the weakest known assumption (even in the linear case) under which a solution in $W^{1,2,p}$ always exists for Dirichlet problems. Moreover, counter-examples in Meyers, (1963) or Ural’ceva, (1967) suggest that this condition is difficult to relax.
@@
The final assumption is used to apply a comparison principle to (1). It is not needed for our existence results, but it is essential to guarantee the uniqueness of a solution to (1).
@@
There exists $\kappa>0$ such that, for all $(t,x)\in\mathcal{Y}$, $q\in\mathbb{R}^{d}$, and $M\in\mathbb{S}^{d}$,
$r\to F(t,x,r,q,M)-\kappa r$
@@
Our main result establishes the existence, uniqueness, and regularity of solutions to (1). [^8] The structure condition (SC) guarantees that any solution of $u_{t}+F[u]=0$ with $u=b$ on $\partial\mathcal{Y}$ is bounded below. Therefore, by choosing $g$ small enough, Theorem 1 also guarantees the existence, uniqueness, and regularity of the solution without obstacles.
@@
Suppose that Assumptions 17 hold for some $p\in(d+2,\infty)$. [^9] The parameter $p$ appears in Assumptions 3 and 7. Then (1) has an $L^{p}$-solution $u$.
Moreover, for all compact $\mathcal{Y}^{\prime}\subset\mathcal{Y}$, there exists $C=C\left(d,p,\lambda,\Lambda,R,diam\left(\mathcal{X}\right),T,dist(\mathcal{Y}^{\prime},\mathcal{Y})\right)\in\mathbb{R}_{+}$ such that
$\left\|u\right\|_{W^{1,2,p}\left(\mathcal{Y^{\prime}}\right)}\leq C\left(1+\underset{a\in A}{\sup}\,\left\|g^{a}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right).$ (2)